shift summary Oct 13,1999 Day

arenton@uvahee.phys.virginia.edu
Wed, 13 Oct 1999 16:48:43 -0400

Shift Summary for day shift, 13/oct/99

On Shift:Mike Arenton, Val Prasad

Hours of beam delivered at all intensities: 5.25
Hours of beam delivered at nominal intensity: 5.0
Hours of E832 data (special runs not included):4.5
Hours of special runs (laser scans, muon runs...):

Accounting of lost time: finish Musep repair, tape change

===============================================================

Runs ended this shift:

Run 14831 - 11.1M events - Beam_799

===============================================================

Tapes brought to FCC this shift:

Accesses during shift: none

===============================================================

Description of Shift:

Beam came back on at 9:45, mostly smooth running thereafter.
daqmon was hung up, so run ended when tapes filled up without
warning.

run spill time comment
----------------------------------------------------
14381 09:45 beam returns, initially at 2E12

10:00 intensity back to 9 E12.

334 10:05 TRD 3 tripped off, notice sweep voltage drawing
current,

344 10:21 TRD 3 back on, reduce sweep voltage to 238V (from 242)

375 11:03 spill very bad, MCR has cleared memory of QXR.

382 spill back to normal.

11:10 notice that V' 1-15 latch is in slice 6. a KFREND
done shortly thereafter to clear Ktevana errors
fixes this.

11:20 twice trigger hung, requiring cycling power on
laser crate repeater.

11:30 sometime about here daqmon stopped updating, only
noticed about an hour later.

11:57 -12.02 beam off SY fridge permit.

492 13:41 ktevana errors - KFREND

14:50 onsplit asks for UPM015 on ktev4

14:53 DDD veto stuck, ktev2 very slow,... finally figure
out that we've reached the end of tapes. Run
wouldn't stop, saved most of L3 histo's

15:20 after fresh start, new run 14382 with new tapes.